OpenStudy (anonymous):

A college student is looking at her monthly checking account records. On September 1, 2008, her checking account held a balance of $1,050. At the end of March 2009, her checking account held a balance of $800. What is the monthly rate of change for the student’s checking account over the period September 1, 2008 through March 31, 2009? help please

OpenStudy (anonymous):

$250

OpenStudy (anonymous):

i thought so . thanks

OpenStudy (anonymous):

wait no thats wrong ...

OpenStudy (jack1):

monthly charge 1st: how many months between Sept 1 2008 and March 31 2009?

OpenStudy (anonymous):

6

OpenStudy (jack1):

yeah, but 7 including sept and march

OpenStudy (jack1):

sept oct nov dec jan feb mar

OpenStudy (jack1):

so initial balance was 1050 and final balance was 800 so amount lost is??

OpenStudy (anonymous):

oh okay

OpenStudy (jack1):

all good ;)

OpenStudy (anonymous):

but when i answered that it said i was wrong

OpenStudy (jack1):

yep, but it wants the answer as a monthly amount so $250 total lost in 7 months = 250/7 per month = about $35.71/month

OpenStudy (jack1):

is that what you got before?

OpenStudy (anonymous):

yes

OpenStudy (jack1):

ok, so the answer isn't linear then 35.71 is based on a fixed rate of change per month so lets try a % based answer

OpenStudy (jack1):

so A = P (1+rt) Amount (A) = Principal (P) ( 1 + Rate(r) x Time (t) ) have u used this formula before?

OpenStudy (jack1):

800 = 1050 (1 + 7r) 800/1050 = (1 + 7r) r = -0.034 per month as a percentage = -3.4% per month

OpenStudy (anonymous):

no i havent. but nevermind. thank you though, i appreciate your kindness ...

OpenStudy (jack1):

k sorry